An object moves along the x-axis with its position x, in meters, given as a function of time t, in seconds, by x(-1.417-9.491+ 4.68 r(t)-1.41 t2-9.4% + 4.68 What is the objects velocity at time t 1.01 s? Number m/ s
Answer #1

Solution:

given x(t) = 1.41 t^2 - 9.49 t + 4.68

=> v(t) = dx(t)/dt

=> v(t) = 2.82 * t - 9.49

=> v(1.01) = 2.82 * 1.01 - 9.49

=> v(1.01) = -6.6418 m/s

please comment for queries.

please rate thanks.